with high levels of Carbon monoxide (Co. F Shifts Night Morn 26 24 62. Influenza Headache 32 Symptoms 29 60 and work shift At 5% signica nce level can you conclude that there is a relationship between the
Answer #1

Given table data is as below

MATRIX col1 col2 col3 TOTALS
row 1 26 13 18 57
row 2 24 32 6 62
row 3 10 16 5 31
TOTALS 60 61 29 N = 150

------------------------------------------------------------------

calculation formula for E table matrix

E-TABLE col1 col2 col3
row 1 row1*col1/N row1*col2/N row1*col3/N
row 2 row2*col1/N row2*col2/N row2*col3/N
row 3 row3*col1/N row3*col2/N row3*col3/N

------------------------------------------------------------------

expected frequecies calculated by applying E - table matrix formulae

E-TABLE col1 col2 col3
row 1 22.8 23.18 11.02
row 2 24.8 25.213 11.987
row 3 12.4 12.607 5.993

------------------------------------------------------------------

calculate chisquare test statistic using given observed frequencies, calculated expected frequencies from above

Oi Ei Oi-Ei (Oi-Ei)^2 (Oi-Ei)^2/Ei
26 22.8 3.2 10.24 0.449
13 23.18 -10.18 103.632 4.471
18 11.02 6.98 48.72 4.421
24 24.8 -0.8 0.64 0.026
32 25.213 6.787 46.063 1.827
6 11.987 -5.987 35.844 2.99
10 12.4 -2.4 5.76 0.465
16 12.607 3.393 11.512 0.913
5 5.993 -0.993 0.986 0.165
chi^2 o = 15.727

------------------------------------------------------------------

set up null vs alternative as

null, Ho: no relation b/w symptoms and the work shift

alternative, H1: exists a relation b/w symptoms and the work shift

level of significance, ? = 0.05

from standard normal table, chi square value at right tailed, chi^2 ?/2 =9.488

since our test is right tailed,reject Ho when chi^2 o > 9.488

we use test statistic chi^2 o = ?(Oi-Ei)^2/Ei

from the table , chi^2 o = 15.727

critical value

the value of |chi^2 ?| at los 0.05 with d.f (r-1)(c-1)= ( 3 -1 ) * ( 3 - 1 ) = 2 * 2 = 4 is 9.488

we got | chi^2| =15.727 & | chi^2 ? | =9.488

make decision

hence value of | chi^2 o | > | chi^2 ?| and here we reject Ho

chi^2 p_value =0.003

ANSWERS

---------------

null, Ho: no relation b/w symptoms and the work shift

alternative, H1: exists a relation b/w symptoms and the work shift

test statistic: 15.727

critical value: 9.488

p-value:0.003

decision: reject Ho

conclude that exists a relation b/w symptoms and the work shift